# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ 5.1 Connecting the inverse rate with existing regularization techniques
We have a characterization of an optimal regularizer for interpolators, the inverse rate, that is distribution-dependent. However, in this section we show how a wide range of existing regularizers are tightly connected to the inverse rate. We do not aim to fully explore these connections, and we do not claim that the new bounds that we derive are better than existing ones. Our focus here is to show how the (inverse) rate function of a model can be used to unify many existing regularization techniques which were previously assumed to be unrelated.
$\ell_{2}$-norm regularization
This regularizer, also known as weight decay, is widely used by the ML community and it is known to improve generalization even in over-parameterized model classes interpolating the training data (Goodfellow et al., 2016). The following result defines a connection between the inverse rate of a model and its norm, under the widely used assumption that the loss function $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ is Lipschitz with respect to the parameters of the model ${\boldsymbol{\theta}}$ (Li and Orabona, 2019).
@@
In many common ML models, the null vector is a model with null variance, that is, $\mathbf{0}\in{\boldsymbol{\Theta}}_{0}$, because the loss of this model is constant for any data sample. For example, in supervised classification problems with $K$ labels, a neural network with null weights has a constant loss equals to $\ln K$. In these cases, the above result shows that by promoting models with small parameter norm we promote models with small inverse rate and, in consequence, the $\ell_{2}$-norm is a proxy to minimize the inverse rate and works as a regularizer. Figure 2 illustrates for our running example how the use of $\ell_{2}$-norm regularization leads to models with small parameter norm and larger rate function. This aligns with the existing evidence in the literature about the use of $\ell_{2}$-norm for successfully regularizing interpolators.
However, in the context of over-parameterized models interpolating the training data, it is well known that the $\ell_{2}$-norm of the parameters of the model does not correlate well with the generalization error (Jiang et al., 2019). That is easy to explain if we consider that the $\ell_{2}$-norm just defines an upper-bound over the inverse rate, which, as shown in Section 4.1 is an optimal proxy for the generalization of interpolators. The gap of this upper-bound is the ultimate responsible of the failure of $\ell_{2}$-norm as a proxy for the generalization error of an interpolator. In fact, an open question in ML is:
Open Question 4 When does the norm of an interpolator correlate with its generalization error?
The following result provides a characterization of a learning setup when this happens.
@@
This result shows that under the exponential family and a large enough number of data samples, the inverse rate is close to the scaled Riemannian norm of the model parameters under the positive definite matrix $\text{Cov}_{\nu}(s({\boldsymbol{y}},{\boldsymbol{x}}))$ that defines the Riemannian metric. Note that, in this case, the Riemannian norm will be an optimal regularizer as Theorem 5.1 applies and the inverse rate is a perfectly tight proxy for the generalization error of interpolators as shown in Section 4.1. The $\ell_{2}$-norm is obtained when the covariance matrix is a diagonal matrix with constant entries, a condition that requires the statistical independence of the components of the sufficient statistics $s(\mathbf{y},\mathbf{x})$ with respect to the data-generating distribution.
@@
Distance from initialization
The distance from initialization is known to be related to the generalization error of an interpolator (Nagarajan and Kolter, 2019b). In fact, it has been successfully used as a regularizer in Hu et al. (2019). In these works, an implicit assumption was that the initial models are produced through one of the well-established initialization schemes used in neural networks, such as Xavier or He initialization (Goodfellow et al., 2016). These parameter initialization schemes provide an initial set of weights which guarantee that the variance of the activations across the different layers remain constant. And, in consequence, the randomly initialized model makes almost constant predictions. If ${\boldsymbol{\theta}}_{0}$ denotes the initial parameter, we should see that $\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}_{0}))$ is very small. Figure 3 (right) shows how this is the case for a randomly intialized Inception model using He initialization.
@@
Furthermore, as happens with the $\ell_{2}$-norm, in the context of over-parameterized interpolators, distance from intialization is not a reliable proxy for the generalization error (Jiang et al., 2019). Again, this is easy to explain if we consider that distance from intialization is just an upper-bound over of the inverse rate. In consequence, an open question in ML is:
Open Question 5 When the distance from intialization of an interpolator perfectly correlates with its generalization error?
We provide again a characterization for the case where the cumulant generating function of the model $J_{{\boldsymbol{\theta}}}(\lambda)$ is a quadratic function or, equivalently, its second-order Taylor approximation around ${\boldsymbol{\theta}}_{0}\in{\boldsymbol{\Theta}}_{0}$ is exact. In this case, as we show in the proof of the following result, the cumulant generation function can be expressed:
@@
where $\text{Cov}_{\nu}(\cdot)$ is the covariance w.r.t. $\nu$ of the gradient of the log-likelihood of each $({\boldsymbol{x}},{\boldsymbol{y}})$ sample. If so, the inverse rate can be expressed as stated next.
@@
When training highly over-parameterized models using gradient descent, it is well known that parameters remain close to their initial values. This phenomenon is called lazy training and is a standard assumption in many theoretical analysis of deep neural networks (Jacot et al., 2018). But when lazy training takes place, the second-order Taylor approximation of the cumulant could become highly accurate. The above proposition then suggests how the Riemannian distance, induced by $\text{Cov}_{\nu}\big{(}\nabla_{{\boldsymbol{\theta}}}\ln p({\boldsymbol{y}}|{\boldsymbol{x}},{\boldsymbol{\theta}}_{0})\big{)}$, would be an optimal proxy for the generalization error of an interpolator in the lazy training regime. This reasoning also helps to understand the limitations in this regard of the Euclidean-distance.
Input-Gradient and Lipschitz Regularization
Input-gradient regularization is a technique used in deep learning to improve the robustness of neural network models (Gouk et al., 2021). This approach focuses on the gradients of the model’s output with respect to its input, which are indicative of how sensitive the model’s predictions are to changes in the input data.
The following result shows how the inverse rate function and the expected norm of the input-gradient of a model are also related. This result is based on the log-sobolev inequalities (Chafaï, 2004) and relies on the assumption that $\nu({\boldsymbol{y}}|{\boldsymbol{x}})$ is deterministic, to simplify gradients with respect to the target variable, and that $\nu({\boldsymbol{x}},{\boldsymbol{y}})$ is a uniformly strictly log-concave density, an assumption satisfy by a wide range of distributions.
@@
${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)\leq\sqrt{\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}}\sqrt{M{\mathbb{E}}_{\nu}\Big{[}\big{\|}\nabla_{x}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\big{\|}_{2}^{2}\Big{]}}\,.$
@@
From the above result, it is straightforward to derive a connection between the generalization of a model and its Lipschitz constant. If a model ${\boldsymbol{\theta}}$ is Lipschitz with respect to the input data ${\boldsymbol{x}}$ with Lipschitz constant denoted $\text{Lip}({\boldsymbol{\theta}})$, then:
@@
The above inequality can be combined with Proposition 5.5 to derive a new upper bound on the inverse of the rate function and on the generalization error of a model. In consequence, models with a small Lipschitz constant, will tend to have a small generalization error. This aligns fully with the existing literature (Neyshabur et al., 2017) and, also, explains the rationale behind regularization techniques that prevent overfitting by controlling the Lipschitz constant (Gouk et al., 2021).
Summary
In this section, we have been able to theoretically characterize an optimal (distribution-dependent) regularizer for interpolators, even for over-parameterized model classes. This theoretical characterization helps to understand why a wide range of commonly used regularization techniques promotes interpolators with a smaller generalization error. We are not aware of previous works able to characterize optimal regularizers for interpolators and establish such a wide range of connections with existing regularization methods.
We have also analyzed, in specific but relevant situations, when some of these common regularization methods could be optimal for interpolators, even in over-parameterized model classes. In this regard, our work complements other related results like Bartlett et al. (2020), where the Euclidean norm was shown to be an optimal regularizer for linear regression interpolators under some specific assumptions.
@@ 6 Invariances
In many machine learning problems, we know that the input data ${\boldsymbol{x}}$ we observed have been transformed, usually as a result of the measuring process. For example, in the case of sensor readings, by random noise due to non-perfect sensor readings or, in the case of images, these may be rotated, blurred, etc, by the image capturing process (Bronstein et al., 2021). In this section, we discuss how widely used techniques in ML like data augmentation (Shorten and Khoshgoftaar, 2019) and invariant architectures (Bronstein et al., 2021) can be jointly explained, using PAC-Chernoff bounds and the rate function, as techniques to improve learning under these transformed inputs.
@@
where ${\boldsymbol{x}}_{0}$ denotes the un-transformed and un-observed input, which is distributed according to $\nu_{0}$, ${\boldsymbol{x}}$ denotes the observed (transformed) input. See Figure 5 for a graph representation.
@@
Under this assumption, the target value of an input is sampled before its transformation, that is, transforming the inputs does not affect its label. In statistical terms, under Assumption 2, we have that the target random variable ${\boldsymbol{y}}$ is conditionally independent of ${\boldsymbol{x}}$ given ${\boldsymbol{x}}_{0}$. This fact perfectly aligns with many common settings, where, for example, rotating, reflecting or cropping an image does not alter the label of the object been represented in it (Bronstein et al., 2021). This assumption is weaker than the standard used in many related works (Chen et al., 2020; Bronstein et al., 2021), where the conditional generating distribution is assumed to be invariant under any transformation. These works assume that $\forall g\in G$, $\nu({\boldsymbol{y}}|{\boldsymbol{x}})=\nu({\boldsymbol{y}}|g({\boldsymbol{x}}))$. This assumption does not hold, for example, when considering rotation over digits, as the true label of the image of a six rotated 180 degrees changes to a nine.
@@
Open Question 6 Why models interpolating transformed data, as described in Assumption 2, have a higher generalization error?
@@
Summarizing, when input-data is transformed (Assumption 2), the input-data provides less information about the target variable (Proposition 6.1). This causes the distribution of the empirical loss $\hat{L}(\cdot,{\boldsymbol{\theta}})$ to be less concentrated and with higher expected loss $L({\boldsymbol{\theta}})$ (as empirically shown in Figure 6 and theoretically argued in Propositions 6.2 and 6.3).
@@ 6.1 Invariant Architectures
Invariant architectures refers to a kind of neural networks that remain unaffected by certain transformations in the input data (Bronstein et al., 2021). In case of images, these transformations might include scaling, rotation, translation, or other geometrical changes.
@@
These invariant models, such as convolutional neural networks (CNNs) (LeCun et al., 1998), are crucial because they can recognize and process patterns irrespective of their position, scale, or orientation. This invariance is particularly important in tasks like image and speech recognition, where the essence of the data needs to be captured regardless of such variations. Invariant architectures have consistently shown superior performance compared to non-invariant models (Goodfellow et al., 2016).
The generalization error of invariant neural networks has been widely studied before. Different generalization bounds have been shown to have complexity terms which tend to shrink when the architecture of the model is invariant (Sokolic et al., 2017; Elesedy, 2022; Behboodi et al., 2022). Although these bounds provide theoretical grounds on why generalization error is reduced with invariant architectures, the problem lies in that they only show that the complexity measures of the respective bounds are reduced, not the generalization error. However, based on the arguments given in Section 1, there is a growing empirical (Zhang et al., 2017) and theoretical evidence (Nagarajan and Kolter, 2019c; Gastpar et al., 2023; Wang et al., 2024) that many of these bounds are non-tight, or, probably, even vacuous, and do not satisfy basic criteria such as the bound decreasing with larger dataset sizes, as specifically happens, for example, with the bound provided by Behboodi et al. (2022). In consequence, to the best of our knowledge, there are not widely accepted, specific generalization bounds that comprehensively captures the effect of the invariance of an architecture on a interpolator’s generalization error, specially in over-parameterized model classes. An open question is then the following,
Open Question 7 Why do (over-parameterized) invariant interpolators generalize better?

### v4

@@ 6.1 Connecting the Inverse Rate with Existing Regularization Techniques
We have shown that the inverse rate function, a distribution-dependent and oracle element, is an optimal regularizer for interpolators. However, in this section we will show how a wide range of existing regularizers are tightly connected to the inverse rate and, in consequence, to our definition of smoothness. We do not aim to fully explore these connections, and we do not claim that the new bounds that we derive are better than existing ones. Our focus here is to show how the (inverse) rate function of a model can be used to unify many existing regularization techniques which were previously assumed to be unrelated.
6.1.1 Norm $\ell_{2}$ Regularization
This regularizer, also known as weight decay, is widely used by the ML community and it is known to improve generalization even in over-parameterized model classes interpolating the training data (Goodfellow et al.,, 2016). The following result defines a connection between the inverse rate of a model and its norm, under the widely used assumption that the loss function $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ is Lipschitz with respect to the parameters of the model ${\boldsymbol{\theta}}$ (Li & Orabona,, 2019).
@@ 6.1.1 Norm $\ell_{2}$ Regularization
In many common ML models, the null vector is a model with null variance, that is, $\mathbf{0}\in{\boldsymbol{\Theta}}_{0}$, because the loss of this model is constant for any data sample. For example, in supervised classification problems with $K$ labels, a neural network with null weights has a constant loss equal to $\ln K$. In these cases, the above result shows that by promoting models with small parameter norm we achieve models with small inverse rate. In consequence, the $\ell_{2}$-norm is a proxy to minimize the inverse rate and works as a regularizer. Figure 2 illustrates for our running example how the use of $\ell_{2}$-norm regularization leads to models with small parameter norm and larger rate function. This aligns with the existing evidence in the literature about the use of $\ell_{2}$-norm for successfully regularizing interpolators.
However, in the context of over-parameterized models interpolating the training data, it is well known that the $\ell_{2}$-norm of the parameters of the model does not correlate well with the generalization error (Jiang et al.,, 2020). That is easy to explain if we consider that the $\ell_{2}$-norm just defines an upper-bound over the inverse rate, which, as shown in Section 4.1 is an optimal proxy for the generalization of interpolators. The gap of this upper-bound is the ultimate responsible of the failure of $\ell_{2}$-norm as a proxy for the generalization error of an interpolator. In fact, an open question in ML is:
Open Question 6 When does the norm of an interpolator correlate with its generalization error?
The following result provides a characterization of a learning setup where these two elements are correlated.
@@
This result shows that under the exponential family and a large enough number of data samples, the inverse rate is close to the scaled Riemannian norm of the model parameters under the positive definite matrix $\text{Cov}_{\nu}(s({\boldsymbol{y}},{\boldsymbol{x}}))$ that defines the Riemannian metric. Note that, in this case, the Riemannian norm will be an optimal regularizer as Theorem 6.1 applies and the inverse rate is a perfectly tight proxy for the generalization error of interpolators as shown in Section 4.1. The $\ell_{2}$-norm is obtained when the covariance matrix is a diagonal matrix with constant entries, a condition that requires the statistical independence of the components of the sufficient statistics $s({\boldsymbol{y}},{\boldsymbol{x}})$ with respect to the data-generating distribution.
@@
6.1.2 Distance From Initialization
The distance from initialization is known to be related to the generalization error of an interpolator (Nagarajan & Kolter,, 2017). In fact, it has been successfully used as a regularizer in Hu et al., (2020). In these works, an implicit assumption was that the initial models are produced through one of the well-established initialization schemes used in neural networks, such as Xavier or He initialization (Goodfellow et al.,, 2016). These parameter initialization schemes provide an initial set of weights which guarantee that the variance of the activations across the different layers remain constant. In consequence, the randomly initialized model makes almost constant predictions. If ${\boldsymbol{\theta}}_{0}$ denotes the initial parameter, we should see that $\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}_{0}))$ is very small. Figure 3 (right) shows how this is the case for a randomly intialized Inception model using He initialization.
@@ 6.1.2 Distance From Initialization
Furthermore, as happens with the $\ell_{2}$-norm, in the context of over-parameterized interpolators, distance from intialization is not a reliable proxy for the generalization error (Jiang et al.,, 2020). Again, this is easy to explain if we consider that distance from intialization is just an upper-bound over of the inverse rate. In consequence, an open question in ML is:
Open Question 7 When the distance from intialization of an interpolator perfectly correlates with its generalization error?
We provide again a characterization for the case where the cumulant generating function of the model $J_{{\boldsymbol{\theta}}}(\lambda)$ is a quadratic function or, equivalently, its second-order Taylor approximation around ${\boldsymbol{\theta}}_{0}\in{\boldsymbol{\Theta}}_{0}$ is exact. In this case, as we show in the proof of Proposition 6.4, the cumulant generating function can be expressed as
@@
where $\text{Cov}_{\nu}(\cdot)$ is the covariance w.r.t. $\nu$ of the gradient of the log-likelihood of each $({\boldsymbol{x}},{\boldsymbol{y}})$ sample. Under this hypothesis, the inverse rate can be expressed as stated next.
@@
When training highly over-parameterized models using gradient descent, it is well known that parameters remain close to their initial values. This phenomenon is called lazy training and is a standard assumption in many theoretical analysis of deep neural networks (Jacot et al.,, 2018). However, when lazy training takes place, the second-order Taylor approximation of the cumulant could become highly accurate. The above proposition then suggests how the Riemannian distance, induced by $\text{Cov}_{\nu}\big{(}\nabla_{{\boldsymbol{\theta}}}\ln p({\boldsymbol{y}}|{\boldsymbol{x}},{\boldsymbol{\theta}}_{0})\big{)}$, would be an optimal proxy for the generalization error of an interpolator in the lazy training regime. This reasoning also helps to understand the limitations in this regard of the Euclidean-distance.
6.1.3 Input-Gradient and Lipschitz Regularization
Input-gradient regularization is a technique used in deep learning to improve the robustness of neural network models (Gouk et al.,, 2021). This approach focuses on the gradients of the model’s output with respect to its input, which are indicative of how sensitive the model’s predictions are to changes in the input data.
The following result shows how the inverse rate function and the expected norm of the input-gradient of a model are also related. This result is based on the log-sobolev inequalities (Chafaï,, 2004) and relies on the assumption that $\nu({\boldsymbol{y}}|{\boldsymbol{x}})$ is deterministic, to simplify gradients with respect to the target variable, and that $\nu({\boldsymbol{x}},{\boldsymbol{y}})$ is a uniformly strictly log-concave density, an assumption satisfied by a wide range of distributions.
@@ 6.1.3 Input-Gradient and Lipschitz Regularization
${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)\leq\sqrt{\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}}\sqrt{M{\mathbb{E}}_{\nu}\Big{[}\big{\|}\nabla_{{\boldsymbol{x}}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\big{\|}_{2}^{2}\Big{]}}\,.$
@@
From the above result, it is straightforward to derive a connection between the generalization of a model and its Lipschitz constant. If a model ${\boldsymbol{\theta}}$ is Lipschitz with respect to the input data ${\boldsymbol{x}}$ with Lipschitz constant denoted $\text{Lip}({\boldsymbol{\theta}})$, then, by definition, the following inequality is verified:
@@
The above inequality can be combined with Proposition 6.5 to derive a new upper bound on the inverse of the rate function and on the generalization error of a model:
${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)\leq\sqrt{\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}}\sqrt{M\text{Lip}({\boldsymbol{\theta}})}\,.$
In consequence, models with a small Lipschitz constant, will tend to have a small generalization error. This aligns fully with the existing literature (Neyshabur et al., 2017b, ) and, also, explains the rationale behind regularization techniques that prevent overfitting by controlling the Lipschitz constant (Gouk et al.,, 2021).
6.1.4 Summary
In this section, we have theoretically characterized an optimal (distribution-dependent) regularizer for interpolators, even for over-parameterized model classes. This theoretical characterization can be used to understand why a wide range of commonly used regularization techniques promotes interpolators with a smaller generalization error. We are not aware of previous works able to characterize optimal regularizers for interpolators and establish such a wide range of connections with existing regularization methods in the machine learning community.
We have also analyzed, in some specific but not less relevant situations, when could some of these common regularization methods be optimal for interpolators; even in over-parameterized model classes. In this regard, our work complements other related theoretical results, like the ones proposed by Bartlett et al., (2020), where the Euclidean norm was shown to be an optimal regularizer for linear regression interpolators under some specific assumptions.
This section allows to draw connections between our definition of smoothness, as outlined in Section 4, and other existing definitions in the literature. We have shown how the smoothest interpolator has near-optimal performance in Proposition 6.1. The clearest connection with existing measures of smoothness arises from the characterization of the Lipschitz constant, as demonstrated by combining Equation (6) and Proposition 6.5. The Lipschitz constant is frequently cited in the literature as a measure of a model’s smoothness (Gouk et al.,, 2021). In this work, we have shown how models with a small Lipschitz constant have a small inverse rate and, in consequence, are smoother according to our own definition.
@@ 7 Invariances
In many machine learning problems, we know that the input data ${\boldsymbol{x}}$ we observed have been transformed, usually as a result of the measuring process. For example, in the case of sensor readings, by random noise due to non-perfect sensors or, in the case of images, these may be rotated, blurred, etc, by the image capturing process (Bronstein et al.,, 2021). In this section, we discuss how widely used techniques in machine learning, like data augmentation (Shorten & Khoshgoftaar,, 2019) and invariant architectures (Bronstein et al.,, 2021), can be jointly explained, using PAC-Chernoff bounds and the rate function, as techniques to improve learning under these transformed inputs.
@@
where ${\boldsymbol{x}}_{0}$ denotes the un-transformed and un-observed input, which is distributed according to $\nu_{0}$, ${\boldsymbol{x}}$ denotes the observed (transformed) input. See Figure 8 (left) for a graph representation.
@@
Under this assumption, the target value of an input is sampled before its transformation, that is, transforming the inputs does not affect its label. In statistical terms, under Assumption 2, we have that the target random variable ${\boldsymbol{y}}$ is conditionally independent of ${\boldsymbol{x}}$ given ${\boldsymbol{x}}_{0}$. This fact perfectly aligns with many common settings, where, for example, rotating, reflecting or cropping an image does not alter the label of the object been represented in it (Bronstein et al.,, 2021). This assumption is weaker than the usual ones used in many related works (Chen et al.,, 2020; Bronstein et al.,, 2021), where the conditional generating distribution is assumed to be invariant under any transformation. These works assume that $\forall g\in G$, $\nu({\boldsymbol{y}}|{\boldsymbol{x}})=\nu({\boldsymbol{y}}|g({\boldsymbol{x}}))$. This assumption does not hold, for example, when considering rotation over digits, as the true label of the image of a six rotated 180 degrees changes to a nine.
@@
Open Question 8 Why models interpolating transformed data, as described in Assumption 2, have a higher generalization error?
@@
Summarizing, when input-data is transformed (Assumption 2), the input-data provides less information about the target variable (Proposition 7.1). This causes the distribution of the empirical loss $\hat{L}(\cdot,{\boldsymbol{\theta}})$ to be less concentrated and with higher expected loss $L({\boldsymbol{\theta}})$ (as empirically shown in Figure 9 and theoretically argued in Propositions 7.2 and 7.3).
@@ 7.1 Invariant Architectures
Invariant architectures refers to a kind of neural networks that remain unaffected by certain transformations in the input data (Bronstein et al.,, 2021). In case of images, these transformations might include scaling, rotation, translation, or other geometrical changes.
@@
These invariant models, such as convolutional neural networks (CNNs) (LeCun et al.,, 1998), are crucial because they can recognize and process patterns irrespective of their position, scale, or orientation. This invariance is particularly important in tasks like image and speech recognition, where the essence of the data needs to be captured regardless of such variations. Invariant architectures have consistently shown superior performance compared to non-invariant models (Goodfellow et al.,, 2016).
The generalization error of invariant neural networks has been widely studied before. Different generalization bounds have been shown to have complexity terms which tend to shrink when the architecture of the model is invariant (Sokolic et al.,, 2017; Elesedy,, 2022; Behboodi et al.,, 2022). Although these bounds provide theoretical grounds on why generalization error is reduced with invariant architectures, the problem lies in that they only show that the complexity measures of the respective bounds are reduced, not the generalization error. However, based on the arguments given in Section 1, there is a growing empirical (Zhang et al.,, 2017) and theoretical evidence (Nagarajan & Kolter, 2019b, ; Gastpar et al.,, 2024; Wang et al.,, 2024) that many of these bounds are non-tight, or, probably, even vacuous, and do not satisfy basic criteria such as the bound decreasing with larger dataset sizes, as specifically happens, for example, with the bound provided by Behboodi et al., (2022). In consequence, to the best of our knowledge, there are not widely accepted, specific generalization bounds that comprehensively captures the effect of the invariance of an architecture on a interpolator’s generalization error, specially in over-parameterized model classes. An open question is then the following,
Open Question 9 Why do (over-parameterized) invariant interpolators generalize better?